### v1

# Finite-sample Identification of Continuous-time Parameter-linear Systems

Abstract
Differentiating noisy, discrete measurements in order to fit an ordinary differential equation can be unreasonably effective. Assuming square-integrable noise and minimal flow regularity, we construct and analyze a finite-difference differentiation filter and a Tikhonov-regularized least squares estimator for the continuous-time parameter-linear system. Combining these contributions in series, we obtain a finite-sample bound on mean absolute error of estimation. As a by-product, we offer a novel analysis of stochastically perturbed Moore-Penrose pseudoinverses.
keywords:
Errors in variables identification, Filtering and smoothing, Continuous Time System Estimation, Nonlinear system identification, Statistical data analysis
thanks: This work was supported by the National Science Foundation CAREER Program (Grant No. 2046292).
This work has been submitted to IFAC for possible publication.
{pf} Proofs omitted in the submitted version appear like this.
1 Introduction
This paper presents an estimator with bounded mean absolute error for the parameter vector $\theta$ of the continuous-time parameter-linear system
$\frac{\partial^{m}y}{\partial t^{m}}$ $=\theta^{\intercal}\phi\left(\frac{\partial^{0}y}{\partial t^{0}},\frac{\partial^{1}y}{\partial t^{1}},\ldots,\frac{\partial^{m-1}y}{\partial t^{m-1}}\right)$ (1)
using noisy, discrete measurements of $y$. Our main result, Thm. 15, establishes that $\operatorname{\mathbb{E}}|\hat{\theta}-\theta|$ can be made arbitrarily small by increasing the amount of data collected.
Continuous-time systems arise from physical constitutive relations such as for Newtonian motion $m\ddot{x}=F$, Fickian diffusion $\dot{q}=\Delta q$, capacitance $C\dot{v}=i$, and the like. These principles can be applied without full knowledge of the underlying physical parameters; after defining the dynamic model, one fits the parameters to output trajectories hoping to recover a physical reality. For example, in a capacitor microphone, sound is transduced to capacitance, which is measured in real-time, e.g. by excitation using a high-frequency electrical signal. This interpretation of system-identification-as-measurement applies to magnetic resonance imaging by Maidens et al. (2016), battery diagnostics by Lin et al. (2019), and so on.
Ljung and Wills (2010) rehearse the necessity and difficulty inherent to continuous-time system identification. There is in general no easy translation from continuous to discrete; to discretize (1) by integrating the dynamics accurately would lose the parameter-linear structure. In light of this, the continuous question is not readily amenable to recent answers in finite-sample identification of discrete-time linear (Tsiamis et al., 2022; Ziemann et al., 2023; Matni and Tu, 2019) and parameter-linear (Mania et al., 2022) systems.
A possible solution to problem (1) is to find $\hat{\theta}$ such that the predicted flow generated by $\hat{\theta}^{\intercal}\phi$ best fits the measured $y$. This is the popular prediction error method (PEM) widely endorsed for nonlinear problems (Ljung, 1998, 2010). Some of the PEM writing refers to classical theory on general nonlinear least squares estimation (Chen et al., 2020; Wu, 1981; Lai, 1994; Jennrich, 1969) whose hypotheses are highly technical, and whose conclusions are mostly asymptotic (as opposed to finite).
Prediction error minimization confers no blanket guarantee, and one can devise a toy instance of (1) to make PEM look bad. Take an unforced first-order linear system over a single time step,
$y(0)$ $=y_{0}\neq 0,$
$y(1)$ $=\alpha y(0);$
observed with i.i.d. Gaussian noise:
$Y_{0}$ $=y(0)+\epsilon_{0},$
$Y_{1}$ $=y(1)+\epsilon_{1},$
$\epsilon_{0},\epsilon_{1}$ $\sim\mathcal{N}(0,1).$
Then the maximum-likelihood estimator $\hat{\alpha}=Y_{1}/Y_{0}$ has a Cauchy distribution scaled by $\alpha$. Even with $n\to\infty$ repetitions, there is no approaching the coveted Cramér-Rao bound found in textbooks such as Bard (1974), as the sampling distribution of $\hat{\alpha}_{n}$ does not even have a mean.
On the other hand, we have finite-sample guarantees when system identification amounts to estimating an operator $\Theta$ in
$Y$ $=\Theta X+\text{noise},$ (2)
where generally, $X$ depends on the initial condition, system input, and observable states; $Y$ is an output; and $\Theta$ is an system representation such as a Hankel matrix or transfer function. After multiplying both sides by a generalized inverse of $X$, estimation error in $\Theta$ is bounded in terms of $X$ and the noise profile (Pillonetto et al., 2022; Tsiamis et al., 2022).
What stops us from reducing (1) to (2)? In the latter, the output is linear in the parameters; in the former, the flow is. All we have to do is reconstruct the unmeasured derivatives of $y$; namely, to cast problem (1) into the form
$Y=\Theta\left(X+\text{error}_{X}\right)-\text{error}_{Y},$ (3)
resembling problem (2), we must differentiate the measurements (van Breugel et al., 2020). Differentiation-based identification, according to conventional wisdom, is hindered by bias (due to high-frequency signal) and fluctuation (due to amplification of high-frequency noise). As an overture to the construction of Section 3, we compare it heuristically to PEM on both counts.
First, the derivative filter introduces bias in the form of generalized aliasing: even without noise, it is never possible to get perfect derivatives for a general smooth $y$, because one cannot with certainty impute what happens in between samples. But suppose, as is often the the case, that the sampling period $\Delta t$ is small compared to the regularity of $y$. When applying PEM to such a case, it would seem viable to integrate the flow using a linear multi-step scheme [^1] A good exposition can be found in Dattani (2008). Butcher (2006) is a comprehensive treatment. with steps aligned to the measurement schedule. The quadrature error is controlled by comparing $y$ to local polynomial approximants. Our differentiator will share the local polynomial approximation, but we manage the error systematically in Lemma 6 and beyond.
Second, differentiation amplifies noise. Indeed, the ideal LTI differentiator has Laplace transfer function $s$. But the noisy derivative is a throwaway intermediate representation that gets smoothed in the next step. For example, to estimate $\dot{y}=\theta$, the time average of $\hat{\dot{y}}$, which is the differentiation-based estimate, is very close to the least-squares PEM estimate.
This paper puts the following contributions in series.
  • An arbitrary-order finite difference filter for differentiating noisy signals: with modest regularity assumptions on trajectory and noise, we are able to bound the mean squared error of the derivative estimates. The filter resembles Savitzky-Golay (Schmid et al., 2022; Niedźwiecki et al., 2021; Ochieng et al., 2023) and orthogonal polynomial filtering (Othmane, 2022; Diekema and Koornwinder, 2012), but emphasizes the bias-variance tradeoff when filtering noisy data, following Sadeghi et al. (2020); Krishnan and Seelamantula (2013).
  • Analysis of the ridge-regression estimator subject to square integrable noise in both the regressors and predictors. Noise in both variables constitutes an errors-in-variables problem familiar in statistics (Gleser et al., 1987; Hirshberg, 2021; Zhang et al., 2020) and system identification (Barbieri and Diversi, 2021; Khorasani and Weyer, 2021; Fosson et al., 2021). Our stochastic pseudoinverse perturbation bound, Lemma 13, reveals the usefulness of Tikhonov regularization in such problems.
1.1 Notation and conventions
[…]
We say that a function is Riemann integrable if its equally-spaced Riemann sums converge.
2 Problem and outline
Our parameter space will be $\Theta=\mathbb{R}^{p}$. The output is $y:[0,T]\times\Theta\to\mathbb{R}$, and the state $\xi$ consists of the derivatives of $y$ of orders 00 through $m-1$. The state dynamics are
$\frac{\partial^{m}}{\partial t^{m}}y(t,\theta)$ $=\theta^{\intercal}\phi(\xi(t,\theta))$ (4)
with a known initial condition $\xi(0,\theta)=z_{0}$.
Assumption 1
The feature map $\phi:\mathbb{R}^{m}\to\mathbb{R}^{p}$ is $\alpha$-Hölder continuous with coefficient $C_{\phi}$.
Assumption 2
$\frac{\partial^{m+1}}{\partial t^{m+1}}y$ exists and is Riemann integrable.
It is possible that the system under identification is closed-loop, and in such a case $u$ is embedded in a non-autonomous $\phi$. In such cases, we do not concern ourselves with choosing the best “$u$”: our tools only speak to how good a particular choice is.
2.1 Stochastic observation and estimation
To represent noisy measurement, we use an abstract probability space consisting of outcomes $\Omega$, $\sigma$-field $\mathcal{F}$, and probability measure $\mathbb{P}$. We write $\operatorname{\mathbb{E}}$ for expectation on this probability space.
We have a zero-mean noise process $w:[0,T]\times\Omega\to\mathbb{R}$ such that for each $t\in[0,T]$, $\omega\mapsto w(t,\omega)$ is an independent random variable. There is no need to negotiate with the fine analytical details of continuous-time random processes, as we never do calculus with $w$; however, our bounds require:
Assumption 3
$\operatorname{\mathbb{E}}w(t)^{2}$ is Riemann integrable in $t$.
We may observe $y$ at $n$ evenly spaced observations through the random variable $z$ with the following law:
$z_{i,n}$ $=y(in^{-1}T,\theta)+w(in^{-1}T),$ (5)
$i$ $=1,2,\ldots,n.$ (6)
Definition 4
A discrete estimator of $\theta$ is a random variable $\hat{\theta}:\Omega\to\Theta$ in the $\sigma$-field generated by $z_{i,n}$ for all $1\leq i\leq n$.
Definition 5
A sequence of discrete estimators $\{\hat{\theta}_{n}\}_{n\in\mathbb{N}}$ is consistent if $\hat{\theta}_{n}\overset{\mathbb{P}}{\longrightarrow}\theta$.
Our problem is: find a consistent sequence of discrete estimators indexed by $n$, the number of samples.
2.2 Plan
In Section 3 we use finite difference filters to get simultaneous derivatives of $y$ of order 00 through $m$ at almost all of the sample times. The only facts carried forward are the error moments eqq. (15) and (17), which we paraphrase as:
$\sum_{i=1}^{n^{\prime}}\sum_{j=0}^{m}\operatorname{MSE}\widehat{\partial_{t}^{j}y(t_{i})}\lesssim\frac{n^{\prime}}{n^{\frac{2}{2m+3}}},$
where $n^{\prime}$ is the length of filtered output (related to subsampling).
Then in Section 4 we plug these into a $\lambda$-ridge regression estimator to get $\hat{\theta}_{n}$. After a lot of real analysis, we will conclude in Cor. 16 that
$\operatorname{\mathbb{E}}\left\|\hat{\theta}_{n}-\theta\right\|\lesssim n^{-\frac{\alpha}{2m+3}},$
in which the dependence on $\sqrt{n^{\prime}}$ has factored completely through terms of the form $\lambda/\sqrt{n^{\prime}}$.
Thus the paper concludes.
2.3 Open question
We wonder whether our classically fashioned Assumption 3 of square-integrability of $w$ allows for nontrivial strengthening of our main result, either absolutely, in the form of an error moment; or relatively as a tail bound stronger than Mean Absolute Deviation $\operatorname{\mathbb{E}}|\hat{\theta}-\operatorname{\mathbb{E}}\hat{\theta}|$. In particular, does $\hat{\theta}$ exhibit concentration as $n^{\prime}\to n$? What about $n^{\prime}\gg n$ if Section 3 is extended to get inter-sample derivatives?
3 Differentiating noisy signals
Use the $n$ random variables $z_{i,n}$ to fashion $n^{\prime}$ evenly-spaced windows, each containing $N$ consecutive samples. The windows can be chosen to be disjoint, $n^{\prime}=n/N$; overlapping, $n^{\prime}=n-N+1$; or anything in between. Intuitively, a higher $n^{\prime}$ is better (why discard data?), but there is a tradeoff when we get to Theorem 15.
Let $n_{i}$ be the first index of the $i$th window. Then the start time of the $i$th window is $t_{i}=n_{i}n^{-1}T$. Let $x^{d}_{i}=\frac{\partial^{d}}{\partial t^{d}}y(t_{i})$, $d=0,1,\ldots,m$. Regard $x_{i}$ as an $\mathbb{R}^{m+1}$ vector, and $\hat{x}_{i}$ likewise (after we construct $\hat{x}_{i}^{d}$). Let $\Delta x_{i}=\hat{x}(t_{i})-x(t_{i})$.
3.1 Building blocks of the filter
Fix $\rho:[0,1]\to(0,\infty)$, a Riemann integrable function. Define an inner product on the polynomials of degree at most [^2] This is a rank-$N$ quadratic form and therefore doesn’t extend to an inner product on general $f$ and $g$. $N-1$:
$\left\langle f,g\right\rangle_{N}$ $=\frac{1}{N}\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)f\left(\frac{j}{N}\right)g\left(\frac{j}{N}\right).$
Let the polynomials $\{p^{d}_{N}:0\leq d\leq N-1\}$, $\deg p^{d}_{N}=d$ be an orthogonal family for each $N$:
$\left\langle p^{d}_{N},p^{d^{\prime}}_{N}\right\rangle_{N}$ $=0\quad\text{when}\quad d\neq d^{\prime}.$
As $\left\langle f,g\right\rangle_{N}$ converges to a Riemann integral as $N\to\infty$, so $p_{N}^{d}$ converges too. For later convenience, define the filter-specific constants
$h_{N}^{d}$ $=\frac{1}{N}\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)p_{N}^{d}\left(\frac{j}{N}\right)\left(\frac{j}{N}\right)^{d}$
$=\left\langle p_{N}^{d},(\cdot)^{d}\right\rangle_{N}$
and
$g_{N}^{d}$ $=\frac{1}{N}\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)\left|p_{N}^{d}\left(\frac{j}{N}\right)\right|$
$=\left\langle p_{N}^{d},\operatorname*{sign}p_{N}^{d}\right\rangle_{N},$
where $\operatorname*{sign}p_{N}^{d}$ is an abuse of notation referring to its alias polynomial of degree at most $N-1$.
For use in analysis, define
$M_{i}^{d}$ $=\sup_{t_{i}\leq t\leq t_{i}+n^{-1}NT}\left|\frac{\partial^{d+1}}{\partial t^{d+1}}y^{(d+1)}(t)\right|$
and
$(s_{i}^{d})^{2}$ $=\frac{1}{N}\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)^{2}p_{N}^{d}\left(\frac{j}{N}\right)^{2}\operatorname{\operatorname{var}}w((n_{i}+j)n^{-1}T).$
3.2 Choice of filter coefficients
Finally, define
$\hat{x}^{d}_{i}$ $=C\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)p^{d}_{N}\left(\frac{j}{N}\right)z_{n_{i}+j,n},$ (7)
where
$C$ $=\frac{d!n^{d}}{N^{d+1}T^{d}h_{N}^{d}}.$
Lemma 6 (Bias of the estimated derivative)
Let $\hat{x}^{d}_{i}$ be as in (7). Then
$\left|\operatorname{\mathbb{E}}\Delta x^{d}_{i}\right|$ $\leq\left(\frac{M_{i}^{d}T}{d+1}\right)\left(\frac{N}{n}\right)\frac{g_{N}^{d}}{h_{N}^{d}}.$
{pf} By Taylor’s theorem, $\operatorname{\mathbb{E}}z_{n_{i}+j,n}$ $=y\left(\frac{(n_{i}+j)T}{n}\right)$ $=\sum_{k=0}^{d}\frac{y^{(k)}(t_{i})}{k!}\left(\frac{NT}{n}\right)^{k}\left(\frac{j}{N}\right)^{k}+R$ (8) where the remainder $R=R(i,d,n,N)$ satisfies $\left|R\right|$ $\leq R_{\text{max}}=\left(\frac{NT}{n}\right)^{d+1}\frac{M_{i}^{d}}{(d+1)!}.$ (9) We now insert this into (7). By the orthogonality property of $p^{d}_{N}$, all of the lower-order Taylor expansion terms of $\operatorname{\mathbb{E}}z_{n_{i}+j,n}$ vanish. $\operatorname{\mathbb{E}}\hat{x}^{d}_{i}$ $=C\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)p_{N}^{d}\left(\frac{j}{N}\right)\frac{x^{d}_{i}}{d!}\left(\frac{NT}{n}\right)^{d}\left(\frac{j}{N}\right)^{d}$ $\phantom{=}+C\sum_{j=1}^{N}\rho\left(\frac{j}{N}\right)p_{N}^{d}\left(\frac{j}{N}\right)R(j,d,n,N)$ (10) By our choice of $C$, the first sum evaluates to $x^{d}_{i}$. $\left|\operatorname{\mathbb{E}}\hat{x}^{d}_{i}-x^{d}_{i}\right|$ $\leq CNR_{\text{max}}g_{N}^{d}$ $=\left(\frac{M_{i}^{d}T}{d+1}\right)\left(\frac{N}{n}\right)\frac{g_{N}^{d}}{h_{N}^{d}}.$
Lemma 7 (Variance of the estimated derivative)
Let $\hat{x}^{d}_{i}$ be as in (7). Then
$\operatorname{\operatorname{var}}\hat{x}^{d}_{i}$ $\leq\frac{n^{2d}}{N^{2d+1}}\left(\frac{d!s_{i,N}^{d}}{T^{d}h_{N}^{d}}\right)^{2}.$
{pf} Direct calculation beginning with $\begin{split}\operatorname{\operatorname{var}}\hat{x}^{d}_{j}=C^{2}\sum_{j=1}^{N}\Bigg{[}&\ \rho\left(\frac{j}{N}\right)^{2}p_{N}^{d}\left(\frac{j}{N}\right)^{2}\\ &\quad\cdot\operatorname{\operatorname{var}}w((n_{i}+j)n^{-1}T)\Bigg{]}.\end{split}$
Theorem 8 (MSE of the estimated derivative)
The mean squared error (MSE) of $\hat{x}_{i}^{d}$ satisfies
$\operatorname{\mathbb{E}}\left\|\Delta x^{d}_{i}\right\|^{2}$ $\leq A_{i,N}^{d}\frac{N^{2}}{n^{2}}+B_{i,N}^{d}\frac{n^{2d}}{N^{2d+1}},$
with constants
[…]
$B_{i,N}^{d}$ $=\left(\frac{d!s_{i,N}^{d}}{T^{d}h_{N}^{d}}\right)^{2}.$
{pf}
Follows from applying the identity $\text{MSE}=\text{Bias}^{2}+\text{Variance}$ to Lemmas 6 and 7.
3.3 Choice of $N$
To pick the filter window size $N$, add up Thm. 8 across all $d=0,1,\ldots,m$.
$\operatorname{\mathbb{E}}\left\|\Delta x_{i}\right\|^{2}$ $\leq\sum_{d=0}^{m}A^{d}_{i,N}\frac{N^{2}}{n^{2}}+\sum_{d=0}^{m}B^{d}_{i,N}\frac{n^{2d}}{N^{2d+1}}$ (11)
We can then bound the total MSE across all windows as
$\operatorname{\mathbb{E}}\sum_{i=1}^{n^{\prime}}\left\|\Delta x_{i}\right\|^{2}$ $\leq\tilde{A}_{N}\frac{N^{2}}{n^{2}}+\tilde{B}_{N}\frac{n^{2m}}{N^{2m+1}},$ (12)
$\tilde{A}_{N}$ $=\sum_{i=1}^{n^{\prime}}\sum_{d=0}^{m}A^{d}_{i,N},$ (13)
$\tilde{B}_{N}$ $=\sum_{i=1}^{n^{\prime}}\sum_{d=0}^{m}B^{d}_{i,N}.$ (14)
Optimizing in $N$, we pick
$N$ $=\left[\frac{(2m+1)\hat{B}_{N}}{2\hat{A}_{N}}\right]^{\frac{1}{2m+3}}n^{\frac{2m+2}{2m+3}},$
so that the total MSE obeys
$\operatorname{\mathbb{E}}\sum_{i=1}^{n^{\prime}}\left\|\Delta x_{i}\right\|^{2}$ $\leq\epsilon_{N}n^{\prime}n^{-\frac{2}{2m+3}},$ (15)
where
$\epsilon_{N}$ $=2(2m+1)^{\frac{1}{2m+3}}\hat{A}_{N}^{\frac{2m+1}{2m+3}}\hat{B}_{N}^{\frac{2}{2m+3}}(n^{\prime})^{-1}$ (16)
tends to a limit as $N\to\infty$.
Likewise, let $\epsilon_{N,m}\leq\epsilon_{N}$ be a constant such that
$\operatorname{\mathbb{E}}\sum_{i=1}^{n^{\prime}}\left\|\Delta x^{m}_{i}\right\|^{2}$ $\leq\epsilon_{N,m}n^{\prime}n^{-\frac{2}{2m+3}}.$ (17)
4 The ridge-regression estimator
4.1 Construction of the estimator
Let $P$ be the projection matrix onto the first $m$ entries. Introduce the shorthand $\phi_{i}=\phi(Px_{i})$, $\hat{\phi}_{i}=\phi(P\hat{x}_{i})$, $\Delta\phi=\hat{\phi}_{i}-\phi_{i}$, and let $\phi$ be Hölder continuous with $\left\|\Delta\phi_{i}\right\|\leq C_{\phi}\left\|\Delta x_{i}\right\|^{\alpha}$. Let $\delta=\left\|\Delta\Phi\right\|$.
Let $\lambda\in(0,\sigma_{\text{min}}(\Phi))$ be a regularization coefficient to be determined later. Let $\Phi\in\mathbb{R}^{n^{\prime}\times p}$ be the matrix having $\phi_{i}^{\intercal},i=1,2,\ldots,n^{\prime}$ as rows, and let $\hat{\Phi}$ be the same, with $\hat{\phi}_{i}$ replacing $\phi_{i}$. Define $R,\hat{R}\in\mathbb{R}^{(p+n^{\prime})\times p}$ as the block matrices
$R$ $=\begin{pmatrix}0_{p\times p}\\ \Phi\end{pmatrix},\quad\hat{R}=\begin{pmatrix}\lambda I_{p}\\ \hat{\Phi}\end{pmatrix}.$ (22)
Define $u,\hat{u}\in\mathbb{R}^{p+n^{\prime}}$ as
$u$ $=\begin{pmatrix}0_{p\times 1}\\ x^{m}_{1}\\ x^{m}_{2}\\ \vdots\\ x^{m}_{n^{\prime}}\end{pmatrix},\quad\hat{u}=\begin{pmatrix}0_{p\times 1}\\ \hat{x}^{m}_{1}\\ \hat{x}^{m}_{2}\\ \hat{\vdots}\\ \hat{x}^{m}_{n^{\prime}}\end{pmatrix}.$ (33)
The system dynamics can be stated as $u=R\theta$. Assuming that $R$ has rank $m$, let $R^{\dagger}$ be the Moore-Penrose pseudoinverse of $R$, leading to the identity
$\theta$ $=R^{\dagger}u$
which motivates the estimator
$\hat{\theta}$ $=\hat{R}^{\dagger}\hat{u}.$
4.2 Error analysis
To analyze the error $\Delta\theta=\hat{\theta}-\theta$, let $v=\hat{u}-u$ and $S=\hat{R}^{\dagger}-R^{\dagger}$.
$\Delta\theta$ $=\left(R^{\dagger}+S\right)\left(u+v\right)-R^{\dagger}u$ (34)
$=R^{\dagger}v+S\left(u+v\right)$ (35)
$\left\|\Delta\theta\right\|$ $\leq\frac{\left\|v\right\|}{\sigma_{\text{min}}(R)}+\left\|S\right\|\left\|u\right\|+\left\|S\right\|\left\|v\right\|$ (36)
To extract the dependence on $n^{\prime}$, we introduce the variables
$\bar{\sigma}$ $=(n^{\prime})^{-1/2}\sigma_{\text{min}}(\Phi),$ (37)
$\bar{U}$ $=(n^{\prime})^{-1/2}\left\|u\right\|$ (38)
These are scaled so that they approach limits as $n\to\infty$.
Lemma 9 (Moment of $v$)
For all $r\in(0,2]$, the random variable $v$ satisfies
$\operatorname{\mathbb{E}}\left\|v\right\|^{r}$ $\leq\frac{\epsilon_{N,m}^{r/2}(n^{\prime})^{r/2}}{n^{\frac{r}{2m+3}}}.$
{pf} By (17), $\operatorname{\mathbb{E}}\left\|v\right\|^{2}$ $=\operatorname{\mathbb{E}}\sum_{i=1}^{n^{\prime}}\left\|\Delta x_{i}^{m}\right\|^{2}$ $\leq\epsilon_{N,m}n^{\prime}n^{-\frac{2}{2m+3}}.$ (39) For all $0<r\leq 2$, Jensen’s gives $\operatorname{\mathbb{E}}\left\|v\right\|^{r}$ $\leq\frac{\epsilon_{N,m}^{r/2}(n^{\prime})^{r/2}}{n^{\frac{r}{2m+3}}}.$ (40)
Lemma 10 (Moment of $\delta$)
The random variable $\delta$ satisfies
$\operatorname{\mathbb{E}}\delta^{2/\alpha}$ $\leq\frac{C_{\phi}^{2/\alpha}\epsilon_{N}(n^{\prime})^{1/\alpha}}{n^{\frac{2}{2m+3}}}.$
{pf} Bounding $\delta$ by the Frobenius norm of $\Delta\Phi$, we have $\delta^{2}$ $\leq\sum_{i=1}^{n^{\prime}}\left\|\Delta\phi_{i}\right\|^{2}$ (41) $\leq C_{\phi}^{2}\sum_{i=1}^{n^{\prime}}\left\|\Delta x_{i}\right\|^{2\alpha}$ (42) By Hölder’s inequality with exponents $1/\alpha$ and $1/(1-\alpha)$, $\leq C_{\phi}^{2}\left(\sum_{i=1}^{n^{\prime}}\left(\left\|\Delta x_{i}\right\|^{2\alpha}\right)^{1/\alpha}\right)^{\alpha}\left(\sum_{i=1}^{n^{\prime}}1^{\frac{1}{1-\alpha}}\right)^{1-\alpha}$ (43) Raising both sides to the power $1/\alpha$ and taking expectations, $\delta^{2/\alpha}$ $\leq C_{\phi}^{2/\alpha}(n^{\prime})^{1/\alpha-1}\sum_{i=1}^{n^{\prime}}\left\|\Delta x_{i}\right\|^{2}$ (44) $\operatorname{\mathbb{E}}\delta^{2/\alpha}$ $\leq C_{\phi}^{2/\alpha}(n^{\prime})^{1/\alpha-1}\sum_{i=1}^{n^{\prime}}\operatorname{\mathbb{E}}\left\|\Delta x_{i}\right\|^{2}$ (45) Applying (15), $\leq\frac{C_{\phi}^{2/\alpha}\epsilon_{N}(n^{\prime})^{1/\alpha}}{n^{\frac{2}{2m+3}}}.$ (46)
To control $S$, we state and then apply a self-contained technical lemma, proven in Appendix A.
Lemma 11
Let $A\in\mathbb{R}^{m\times n}$ have full column rank and minimal singular value $\sigma$, and let $D$ be a real matrix of the same dimensions as $A$, having operator norm $\delta$. For any $\lambda>0$, let
[…]
$S$ $=B-\hat{B}.$
Then
$\left\|S\right\|$ $\leq\frac{2(\lambda+\delta)}{\sigma(\lambda+(\sigma-\delta)^{+})}.$
Lemma 12 (High moment of $S$)
The random matrix $S=\hat{R}^{\dagger}-R^{\dagger}$ satisfies
$\operatorname{\mathbb{E}}\left\|S\right\|^{2/\alpha}$ $\leq\frac{2^{4/\alpha-1}\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}+\frac{2^{4/\alpha-1}}{\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}.$
{pf} Applying Lemma 11 to the facts of the case, $\left\|S\right\|$ $\leq\frac{{2}\left(\delta+\lambda\right)}{\overline{\sigma}\sqrt{n^{\prime}}\left((\overline{\sigma}\sqrt{n^{\prime}}-\delta)^{+}+\lambda\right)}$ (47) As we will see that $\delta\overset{\mathbb{P}}{\longrightarrow}0$ as $n\to\infty$, the following bound is very conservative for large $n$, but it is difficult to do better when we need a high moment of $S$. $\leq\frac{{2}\left(\delta+\lambda\right)}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}$ (48) Using the fact that $(a+b)^{p}\leq 2^{p-1}(a^{p}+b^{p})$ for $a,b\geq 0$ and $p\geq 1$, $\left\|S\right\|^{2/\alpha}$ $\leq 2^{4/\alpha-1}\frac{\delta^{2/\alpha}+\lambda^{2/\alpha}}{\lambda^{2/\alpha}\sigma_{\text{min}}^{2/\alpha}(\Phi)}$ (49) $\operatorname{\mathbb{E}}\left\|S\right\|^{2/\alpha}$ $\leq\frac{2^{4/\alpha-1}\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}+\frac{2^{4/\alpha-1}}{\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}$ (50)
This bound is not sharp enough to control the $Su$ error term by Hölder’s inequality. Some improvement is possible by dividing and conquering $S=S\bm{1}_{\delta\leq\lambda}+S\bm{1}_{\delta>\lambda}$, head and tail. Restricted to the head, $S$ satisfies, in the terms of Lemma 11, $\left\|S\right\|\lesssim\frac{\lambda}{\sigma^{2}}$; restricted to the tail $S$ satisfies the a priori $S\leq\frac{1}{\sigma}$, which can be combined by Hölder’s inequality with the decaying $\operatorname{\mathbb{E}}\bm{}{\delta>\lambda}$. However since $\lambda<\sigma$ by hypothesis and we will later prefer $\lambda\ll\sigma$, it behooves to get tighter control of $S$ on the event $\left\{\lambda<\delta\leq\sigma\right\}$.
We first state a self-contained technical lemma, proved in Appendix B, concerning Tikhonov pseudoinverses of perturbed matrices.
Lemma 13 (Head-body-tail)
Let $A\in\mathbb{R}^{m\times n}$ have full column rank and minimal singular value $\sigma$, and let $D$ be a real random matrix of the same dimensions as $A$, having operator norm $\delta$. For $\lambda\in(0,\sigma)$, let
[…]
$S$ $=B-\hat{B}.$
Then
[…]
where
[…]
and
[…]
Now we return the problem at hand by going through the steps of Lemma 13, in which we have an integrability condition on $\delta$.
Lemma 14
The random matrix $\left\|S\right\|$ satisfies
[…]
where
[…]
and
[…]
{pf} We can immediately transcribe $s_{\text{head}}$ from the statement of Lemma 13. We proceed with $s_{\text{body}}$, using Chebyshev’s inequality to bound $\operatorname{\mathbb{P}}(\delta>v)$. Using $\sigma=\overline{\sigma}\sqrt{n^{\prime}}$ for temporary brevity, $3\int_{\lambda}^{\sigma}\frac{(\lambda+\sigma)\operatorname{\mathbb{P}}(\delta>v)}{\sigma(\lambda+\sigma-v)^{2}}\,\operatorname{d}v$ $\leq 3\operatorname{\mathbb{E}}\delta^{2}\int_{\lambda}^{\sigma}\frac{\lambda+\sigma}{(\lambda+\sigma-v)^{2}v^{2}\,\operatorname{d}v}$ $=\frac{6\operatorname{\mathbb{E}}\delta^{2}}{\sigma}\left[\frac{2\log(\sigma/\lambda)}{(\sigma+\lambda)^{2}}+\frac{\sigma-\lambda}{(\sigma+\lambda)\sigma\lambda}\right]$ “Rounding up” for simplicity, $\leq\frac{6\operatorname{\mathbb{E}}\delta^{2}}{\sigma}\left[\frac{2\log(\sigma/\lambda)}{\sigma^{2}}+\frac{1}{\sigma\lambda}\right]$ Applying Lemma 10 we get this lemma’s $s_{\text{body}}$. $\leq\frac{6C_{\phi}^{2}\epsilon_{N}^{\alpha}}{n^{\frac{2\alpha}{2m+3}}}\left[\frac{2\log(\overline{\sigma}\sqrt{n^{\prime}}/\lambda)}{\overline{\sigma}^{3}\sqrt{n^{\prime}}}+\frac{1}{\lambda\overline{\sigma}^{2}}\right].$ As for $s_{\text{tail}}$, by Hölder’s inequality, $\operatorname{\mathbb{E}}\frac{4\delta}{\sigma\lambda}\bm{1}_{\delta>\sigma}$ $\leq\frac{4}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}\left(\operatorname{\mathbb{E}}\delta^{2/\alpha}\right)^{\alpha/2}\left[\operatorname{\mathbb{P}}\left(\delta>\overline{\sigma}\sqrt{n^{\prime}}\right)\right]^{(2-\alpha)/2}$ $\leq\frac{4}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}\left(\operatorname{\mathbb{E}}\delta^{2/\alpha}\right)^{\alpha/2}\left[\frac{\operatorname{\mathbb{E}}\delta^{2/\alpha}}{(\overline{\sigma}^{2}n^{\prime})^{1/\alpha}}\right]^{(2-\alpha)/2}$ $=\frac{4}{\lambda\overline{\sigma}\sqrt{n^{\prime}}}\frac{\operatorname{\mathbb{E}}\delta^{2/\alpha}}{(\overline{\sigma}^{2}n^{\prime})^{\frac{2-\alpha}{2\alpha}}}$ $=\frac{4\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}$ $\leq\frac{4C_{\phi}^{2/\alpha}\epsilon_{N}}{\lambda\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}}.$
4.3 Bound on $\Delta\theta$
We have now controlled each term in (36),
$\left\|\Delta\theta\right\|$ $\leq\frac{\left\|v\right\|}{\sigma_{\text{min}}(R)}+\left\|S\right\|\left\|u\right\|+\left\|S\right\|\left\|v\right\|.$
Theorem 15 (First moment of $\Delta\theta$)
The random vector $\Delta\theta$ satisfies
$\begin{split}\operatorname{\mathbb{E}}\left\|\Delta\theta\right\|&\leq\frac{\epsilon_{N,m}^{1/2}}{\overline{\sigma}n^{\frac{1}{2m+3}}}+\frac{4\bar{U}\lambda}{\overline{\sigma}^{2}\sqrt{n^{\prime}}}\\ &\quad+\frac{6C_{\phi}^{2\alpha}\bar{U}\epsilon_{N}^{\alpha}}{n^{\frac{2\alpha}{2m+3}}}\left[\frac{2\log(\overline{\sigma}\sqrt{n^{\prime}}/\lambda)}{\overline{\sigma}^{3}}+\frac{\sqrt{n^{\prime}}}{\lambda\overline{\sigma}^{2}}\right]\\ &\quad+\frac{4C_{\phi}^{2}\bar{U}\epsilon_{N}\sqrt{n^{\prime}}}{\lambda\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}}\\ &\quad+\frac{2^{2-\frac{\alpha}{2}}\epsilon_{N,m}^{1/2}}{n^{\frac{1}{2m+3}}}\left(\frac{C_{\phi}^{2}\epsilon_{N}(n^{\prime})^{1/\alpha}}{\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}\lambda^{2/\alpha}}+\frac{1}{\overline{\sigma}^{2/\alpha}}\right)^{\alpha/2}.\end{split}$
{pf} According to Lemma 9, $\operatorname{\mathbb{E}}\frac{\left\|v\right\|}{\sigma_{\text{min}}(R)}$ $\leq\frac{\epsilon_{N,m}^{1/2}}{\overline{\sigma}n^{\frac{1}{2m+3}}}$ According to Lemma 14, we can bound $\left\|u\right\|\operatorname{\mathbb{E}}\left\|S\right\|$ as the sum of $\left\|u\right\|s_{\text{head}}$ $=\frac{4\bar{U}\lambda}{\overline{\sigma}^{2}\sqrt{n^{\prime}}},$ $\left\|u\right\|s_{\text{body}}$ $=\frac{6C_{\phi}^{2\alpha}\bar{U}\epsilon_{N}^{\alpha}}{n^{\frac{2\alpha}{2m+3}}}\left[\frac{2\log(\overline{\sigma}\sqrt{n^{\prime}}/\lambda)}{\overline{\sigma}^{3}}+\frac{\sqrt{n^{\prime}}}{\lambda\overline{\sigma}^{2}}\right],$ and $\left\|u\right\|s_{\text{tail}}$ $=\frac{4C_{\phi}^{2}\bar{U}\epsilon_{N}\sqrt{n^{\prime}}}{\lambda\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}}.$ Using Lemmas 12 and 9 in Hölder’s inequality, $\operatorname{\mathbb{E}}\left\|S\right\|\left\|v\right\|$ $\leq\left(\operatorname{\mathbb{E}}\left\|S\right\|^{2/\alpha}\right)^{\alpha/2}\left(\operatorname{\mathbb{E}}\left\|v\right\|^{2/(2-\alpha)}\right)^{(2-\alpha)/2}$ $\leq\left(\frac{2^{4/\alpha-1}\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}+\frac{2^{4/\alpha-1}}{\overline{\sigma}^{2/\alpha}(n^{\prime})^{1/\alpha}}\right)^{\alpha/2}$ $\quad\cdot\left(\frac{\epsilon_{N,m}^{1/2}(n^{\prime})^{1/2}}{n^{\frac{1}{2m+3}}}\right)$ $\leq\frac{2^{2-\frac{\alpha}{2}}\epsilon_{N,m}^{1/2}}{n^{\frac{1}{2m+3}}}\left(\frac{\operatorname{\mathbb{E}}\delta^{2/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}}+\frac{1}{\overline{\sigma}^{2/\alpha}}\right)^{\alpha/2}$ $\leq\frac{2^{2-\frac{\alpha}{2}}\epsilon_{N,m}^{1/2}}{n^{\frac{1}{2m+3}}}\left(\frac{C_{\phi}^{2/\alpha}\epsilon_{N}(n^{\prime})^{1/\alpha}}{\lambda^{2/\alpha}\overline{\sigma}^{2/\alpha}n^{\frac{2}{2m+3}}}+\frac{1}{\overline{\sigma}^{2/\alpha}}\right)^{\alpha/2}.$
The terms in this bound go both ways with respect to $n^{\prime}$. On one hand, there is $(n^{\prime})^{-1/2}$ scaling reminiscent of the concentration of measure in conventional linear regression. On the other hand, pseudo-inverting a random matrix with $n^{\prime}$ rows offers $n^{\prime}$ ways to go wrong, and we get a $(n^{\prime})^{1/2}$ scaling in those terms.
A parsimonious choice of $\lambda$ is
$\frac{\lambda^{*}}{\sqrt{n^{\prime}}}$ $=\frac{C_{\phi}\epsilon_{N}^{\alpha/2}}{n^{\alpha/(2m+3)}},$ (51)
which consolidates the two terms of Lemma 12, making the $\left\|S\right\|\left\|v\right\|$ estimate proportional to the $\left\|v\right\|/\sigma_{\text{min}}(R)$ estimate. We apply this choice to Thm. 15 to get the following.
Corollary 16
Under the choice $\lambda=\lambda^{*}$ of (51), $\operatorname{\mathbb{E}}\left\|\Delta\theta\right\|=O(n^{-\alpha/(2m+3)})$ as $n\to\infty$.
Appendix A Proof of Lemma 11
Using Thm. 4.1 of Wedin (1973) for the full-rank perturbation of a full-rank rectangular matrix,
$\left\|S\right\|$ $\leq\sqrt{2}\left\|\begin{pmatrix}0_{n\times n}\\ A\end{pmatrix}^{\dagger}\right\|\left\|\begin{pmatrix}\lambda I_{n}\\ A+D\end{pmatrix}^{\dagger}\right\|\left\|\begin{pmatrix}\lambda I_{n}\\ D\end{pmatrix}\right\|.$ (58)
Recall that the operator norm of a matrix’s pseudoinverse is the inverse of its least singular value. Thus,
$\left\|\begin{pmatrix}0_{n\times n}\\ A\end{pmatrix}^{\dagger}\right\|$ $=\frac{1}{\sigma}$ (61)
and
$\left\|\begin{pmatrix}\lambda I_{n}\\ A+D\end{pmatrix}^{\dagger}\right\|^{-2}$ $=\lambda^{2}+\sigma_{\text{min}}^{2}(A+D).$ (64)
By Jensen’s and Weyl’s inequalities,
$\left\|\begin{pmatrix}\lambda I_{n}\\ A+D\end{pmatrix}^{\dagger}\right\|^{-1}$ $\geq\frac{\lambda+\sigma_{\text{min}}(A+D)}{\sqrt{2}}.$ (67)
$\geq\frac{\lambda+(\sigma-\delta)^{+}}{\sqrt{2}}$ (68)
where $\left(\cdot\right)^{+}=\max(0,\cdot)$. We bound the last factor of (58) using the triangle inequality.
Appendix B Proof of Lemma 13
Recall the a priori bound from Lemma 11,
$\left\|S\right\|$ $\leq\frac{2(\delta+\lambda)}{\sigma(\lambda+(\sigma-\delta)^{+})}.$ (69)
We will integrate $\operatorname{\mathbb{E}}\left\|S\right\|$ over three zones of $\delta$: “head,” $[0,\lambda]$; “body,” $(\lambda,\sigma]$; and “tail,” $(\sigma,\infty)$.
First, we have the deterministic
$\left\|S\right\|\bm{1}_{\delta\leq\lambda}$ $\leq\frac{4\lambda}{\sigma^{2}}\bm{1}_{\delta\leq\lambda}$ (70)
Second, define the function $F(x)=\frac{2(\lambda+x)}{\sigma(\lambda+\sigma-x)}$. We will execute a “layer cake” rearrangement of $F(\delta)$.
$\left\|S\right\|\bm{1}_{\lambda<\delta\leq\sigma}$ $\leq F(\delta)\bm{1}_{\lambda<\delta\leq\sigma}$
By the Fundamental Theorem of Calculus,
$=\left(F(\lambda)+\int_{\lambda}^{\delta}F^{\prime}(v)\,\operatorname{d}v\right)\bm{1}_{\lambda<\delta\leq\sigma}$ (71)
$=F(\lambda)\bm{1}_{\lambda<\delta}+\left(\int_{\lambda}^{\delta}F^{\prime}(v)\,\operatorname{d}v\right)\bm{1}_{\lambda<\delta\leq\sigma}$ (72)
Let us combine the former term of (72) with (70), and call it $s_{\text{head}}$. The latter term we will call $s_{\text{body},0}$. We get the bound
$\left\|S\right\|\bm{1}_{\delta\leq\sigma}$ $\leq s_{\text{head}}+s_{\text{body},0},$
$s_{\text{body},0}$ $=\left(\int_{\lambda}^{\delta}F^{\prime}(v)\,\operatorname{d}v\right)\bm{1}_{\lambda<\delta\leq\sigma}$
$=\left(\int_{0}^{\infty}F^{\prime}(v)\bm{1}_{\lambda<v<\delta}\,\operatorname{d}v\right)\bm{1}_{\lambda<\delta\leq\sigma}$
We conservatively (giving up the predicate $\delta\leq\sigma$) rework the logic so that the bounds of integration do not depend on $\delta$.
$\leq\int_{0}^{\infty}F^{\prime}(v)\bm{1}_{\lambda<v<\sigma}\bm{1}_{v<\delta}\,\operatorname{d}v$
$=\int_{\lambda}^{\sigma}F^{\prime}(v)\bm{1}_{v<\delta}\,\operatorname{d}v$
Taking the expectation of both sides and applying Tonelli’s theorem,
$\operatorname{\mathbb{E}}s_{\text{body},0}$ $\leq\int_{\lambda}^{\sigma}F^{\prime}(v)\operatorname{\mathbb{P}}(\delta>v)\,\operatorname{d}v$
By Chebyshev’s inequality,
$=\int_{\lambda}^{\sigma}\frac{2(2\lambda+\sigma)\operatorname{\mathbb{P}}(\delta>v)}{\sigma(\lambda+\sigma-v)^{2}}\,\operatorname{d}v$
To balance $\lambda$ and $\sigma$, we trade in $\lambda/2$ for $\sigma/2$.
$\leq 3\int_{\lambda}^{\sigma}\frac{(\lambda+\sigma)\operatorname{\mathbb{P}}(\delta>v)}{\sigma(\lambda+\sigma-v)^{2}}\,\operatorname{d}v$
We call this last quantity $s_{\text{body}}$. When $\delta$ is large, (69) becomes
$\left\|S\right\|\bm{1}_{\delta>\sigma}$ $\leq\frac{3(\delta+\lambda)}{\sigma\lambda}\bm{1}_{\delta>\sigma}.$
We do not lose much by simplifying $\lambda<\sigma<\delta$.
$\leq\frac{4\delta}{\sigma\lambda}\bm{1}_{\delta>\sigma}.$
The expectation of this quantity constitutes $s_{\text{tail}}$.